will:

The sum of two integers is at most -57. One integer is 23. What is the other integer? Use x to represent the variable.

Vocaloid:

hint: "at most" means "less than or equal to" so x + 23 <= -57 so solve for x

will:

how i dont know how to do this im helping a friend and i just need an answer im not will im kylie

Vocaloid:

subtract 23 from both sides

Vocaloid:

x + 23 - 23 = ?

Vocaloid:

hint: the + 23 and the - 23 cancel each other out... so what would be left?

will:

x

will:

x=0

Vocaloid:

very nice! so we have: x <= -57 - 23 that means that x <= -80 so you can pick any integer that is -80 or less i'm not sure why it asks for only one value of x though

Vocaloid:

**no, x is not equal to 0

will:

omg i dont understand can you explain better

Vocaloid:

sure let's go back to the beginning: we have some unknown integer "x". we know that "the sum of the integer and 23 is equal OR less than -57" that gives us: x + 23 <= -57

Vocaloid:

now, we have "x + 23" on the left side. we want to "isolate" x. this means we want x by itself, so we must get rid of that 23 somehow. the opposite of + 23 is - 23, so we can subtract to get: x + 23 - 23 = x

Vocaloid:

however, since we subtracted 23 from the left side, we must also subtract 23 from the right side - 57 - 23 = -80 so x <= -80 this means that x must be an integer that is less than or equal to -80. so you can pick -80, or -81, or -82... etc.

will:

dont worry its will again she took my computer

Vocaloid:

ok do you understand it a bit better now or are you still confused?

will:

i understand she dont know how to do algabra